# RoCoFT: Efficient Finetuning of Large Language Models with Row-Column Updates

### v2

@@ Abstract
We propose RoCoFT, a parameter-efficient fine-tuning method for large-scale language models (LMs) based on updating only a few rows and columns of the weight matrices in transformers. Through extensive experiments with medium size LMs like BERT and RoBERTa, and larger LMs like Bloom-7B, Llama2-7B and Llama2-13B, we show that our method gives comparable or better accuracies than state-of-art PEFT methods while also being more memory and computation-efficient. We also study the reason behind the effectiveness of our method with tools from neural tangent kernel theory. We empirically demonstrate that our kernel, constructed using a restricted set of row and column parameters, are numerically close to the full-parameter kernel and gives comparable classification performance. Ablation studies are conducted to investigate the impact of different algorithmic choices, including the selection strategy for rows and columns as well as the optimal rank for effective implementation of our method. Code is available at: https://github.com/Kowsher/RoCoFT.
@@ 1 Introduction
Adapting Large Language Models (LLMs) to different downstream applications is the current prevailing paradiagm for solving many Natural Language Processing (NLP) tasks, such as sentiment analysis, machine translation, question answering, named entity recognition, and text summarization. Large language models like GPT-4 (Achiam et al., 2023) and Llama (Touvron et al., 2023) are trained on massive amount of text data and contain billions of parameters. They give state-of-art performance on many NLP, mathematical reasoning (Hendrycks et al., 2021; Cobbe et al., 2021), and programming benchmarks(Jiang et al., 2024). Early works on transfer learning with pretrained LLMs, such as BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), use full fine-tuning, which updates all the parameters of the LLMs when adapting to downstream tasks. This approach becomes impractical as language models continue to scale up (Hoffmann et al., 2022), since a separate copy of the model parameters needs to be stored for each downstream application. Updating all the parameters is also prone to overfitting and the loss of LLM capabilities due to catastrophic forgetting (Kirkpatrick et al., 2017), where the model loses previously learned knowledge while adapting to new tasks.
Adaptor methods (Houlsby et al., 2019; Kowsher et al., 2024) resolve this problem of finetuning LLMs by introducing extra modules called adaptors with a small set of independent parameters. Only the parameters in the adaptors need to be optimized during finetuning, and their small size makes it efficient to adapt an LLM to many different tasks. Parameter-Efficient Finetuning (PEFT) is the study of adapting LLMs to downstream applications by finetuning only a very small set of parameters. Many PEFT methods have been proposed, including the popular LoRA (Hu et al., 2021) and its variants (Zhang et al., 2023b; Edalati et al., 2022; Hyeon-Woo et al., 2021), prefix and prompt tuning (Li & Liang, 2021; Lester et al., 2021), and many other more advanced and complex adaptor methods (He et al., 2021; Zeng et al., 2023). These PEFT methods are effective in reducing the number of parameters required to adapt to downstream tasks, while maintaining performance close to full finetuning.
@@
Despite the numerous PEFT methods available, we pose a critical question: can we design even simpler PEFT methods capable of adapting LLMs to diverse downstream tasks in a more efficient way? A simpler method could not only enhance computational and storage efficiency but also offer deeper insights into why PEFT methods succeed as simpler methods are easier to analyze. We answer this question by presenting a new method called RoCoFT, where the LLMs can be efficiently adapted by updating only a small subset of rows or columns in the transformer block weight matrices. We evaluate the effectiveness of our approach across several benchmarks on different language models. Our experimental results demonstrate that RoCoFT outperforms current PEFT techniques in accuracies, requires fewer trainable parameters and has faster training times. We further analyze our method using Neural Tangent Kernel (NTK) theory (Jacot et al., 2018; Malladi et al., 2023), demonstrating that, for a pretrained LLM, the NTKs derived from a restricted set of rows and columns closely resemble those computed from the full parameter set. This substantiates the effectiveness of our proposed method and further suggests that most of the critical features for fine-tuning are already acquired during the pretraining phase. This insight helps explain why many PEFT methods are effective with only so few parameters, as minimal additional learning is required when a strong set of foundational features is already established. The contributions of this paper are:
$\bullet$ We introduce a new PEFT method called RoCoFT which gives comparable or better accuracies than state-of-art PEFT methods, while being more efficient in terms of memory and time complexity. These claims are validated through extensive experiments on language models of different sizes and many benchmark datasets.
@@
$\bullet$ We analyze our method with empirical neural tangent kernels and show that these kernels are close to NTKs defined on the full parameter set, and they give comparable accuracies on many tasks when trained with kernel logistic regression. This explains why our method has performance close to full finetuning from the view of kernel methods.
$\bullet$ We perform extensive ablation studies on the design choices such as which and how many rows and columns to select to facilitate the implementation of our method
@@ 2 Related Works
PEFT Methods: Parameter-Efficient Finetuning (PEFT) methods aim to finetune only a small number of existing or extra parameters of the LLM to achieve results comparable to finetuning all the parameters. Recently, numerous PEFT approaches have been proposed to advance this strategy. LoRA (Hu et al., 2021) and related methods (Zhang et al., 2023b; Kopiczko et al., 2023; Dettmers et al., 2024) modify existing weight matrices of the model by introducing trainable low-rank decomposition matrices, as adapters, into each layer of the Transformer (Vaswani et al., 2017) architecture. With ranks as low as 4 or 8 for many tasks, they significantly reduce both memory usage and computational time. IA3 (Liu et al., 2022) is another adaptor method that only trains scaling vectors for the key, value, and feed-forward weight matrices in the attention mechanism for task adaptation. Prefix-Tuning (Li & Liang, 2021) and Prompt-Tuning (Lester et al., 2021; Kowsher et al., 2023) work by adding task-specific continuous vectors as contexts for inputs and only updates those parameters while keeping the original LLM parameters frozen. MAM adaptors (He et al., 2021) generalize from both LoRA and prefix-tuning under a unified framework. Our method is closer to LoRA and IA3 in that we modify the weight matrices in the transformer architecture. However, unlike these approaches, we introduce no extra parameters and modify the existing parameters in place. BitFit (Zaken et al., 2021) and LayerNorm Tuning (Zhao et al., 2023) finetune only the bias parameters and layernorm parameters respectively and are extremely parameter-efficient. However, unlike LoRA and our method they cannot increase the capacity of the finetuning model by increasing the rank, since the number of bias and layernorm parameters are fixed in a model.
@@
Neural Tangent Kernels:  Jacot et al. (2018) and related studies (Lee et al., 2019) show that the training dynamics of an infinite-width multi-layer neural network with suitable Gaussian initialization can be completely described by a fixed kernel called the Neural Tangent Kernel (NTK). This result is further expanded in Yang (2020) to any architecture for which forward and backpropagation can be expressed via nonlinearities and matrix multiplications. In particular, it it shown that in the infinite-width limit for any modern architecture including models with attention layers, the NTK converges almost surely to a deterministic limit. Although these results are asymptotic, this interesting connection between deep neural networks and kernel methods allows many questions about neural networks to be studied via kernels. For example, Wei et al. (2022) studied the generalization error of representations learned by deep neural networks through kernel regression with NTKs. Recently Malladi et al. (2023) proposed to study the effect of finetuning LLMs through their NTKs, and provided theoretical support to their approach. In this paper, we continue along this line of work to use NTKs to analyze PEFT methods.
@@
Figure 1: A simplified overview of various PEFT methods and RoCoFT (snowflake icon indicates frozen parameters while fire icon indicates trainable parameters)
@@ 3 RoCoFT
PEFT is a collection of methods for transferring pretrained LLMs to downstream tasks by optimizing only a very small set of (additional) parameters. Since most modern LLMs are based on the transformer architecture (Vaswani et al., 2017), there is a line of work in PEFT focusing on modifying the transformer block by freezing most of its parameters and training only a few additional parameters. The method proposed in Houlsby et al. (2019) adds adaptive layers to the transformer blocks, and only parameters in those adaptive layers need to be trained for effective transfer learning. LoRA (Hu et al., 2021) removes the need for adding adaptive layers by directly modifying the weight matrices used in the transformer blocks. There are multiple linear weight matrices in the transformer block taking up most of the parameters, including $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ for the query, key and value matrices in the attention mechanism, and also the weights $\mathbf{W}_{ff}$ for the MLP projection layers. LoRA makes use of a low-rank modification of these weight matrices:
$\mathbf{W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A},$
where $\mathbf{W}_{0}$ is the pretrained weight matrix, $\mathbf{B}$ and $\mathbf{A}$ are low rank matrices of rank $r$, and $\mathbf{W}$ is the weight matrix after finetuning. In this formulation only $\mathbf{B}$ and $\mathbf{A}$ are updated. If $\mathbf{W}$ is of dimensions $d\times k$, $\mathbf{B}$ and $\mathbf{A}$ will be of dimensions $d\times r$ and $r\times k$ respectively. If $r\ll d,k$, this can lead to signficant savings in terms of mermory and runtime. IA3 in (Liu et al., 2022) also modifies the weight matrices in the tranformer block, but instead of a low-rank modification they rescale the key and value matrices, and also the MLP layers using vectors $l_{k}$, $l_{v}$ and $l_{ff}$.
The success of these PEFT methods leads us to ask if there are even simpler methods for modifying the transformer block for effective finetuning. Here we propose modifying only a few rows or columns in the weight matrices of the transformer block, for query, key, value weight matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ and also the weight matrices in the feedforward layer $\mathbf{W}_{ff}$. These can be expressed as:
$\mathbf{W}=\mathbf{W}_{0}+\mathbf{R}\qquad\mbox{and}\qquad\mathbf{W}=\mathbf{W}_{0}+\mathbf{C}$
@@
Intuition on why the method works: One might wonder why such a simple update scheme can be effective for finetuning LLMs for different tasks, and our experiments in the following sections show even updating 1 row or 1 column per weight matrix can be extremely effective. We believe this is related to the phenomenon that most of the knowledge of the LLMs are learned during the pretraining stage, and only very limited learning or adaptation is required during finetuning, as observed in the work Zhou et al. (2024) on the limited number of examples required for finetuning/alignment. In the Appendix we conduct extra experiments to show that even updating a set of randomly selected entries of the weight matrices can work very well. So there is perhaps nothing special about using low-rank or row and column-based matrices. As long as there are sufficient number of parameters spread throughout the LLM for updates, finetuning can be successful. We also give support to this argument of most knowledge already acquired during pretraining by making use of the recent theory of NTK for finetuning LLMs (Malladi et al., 2023; Jacot et al., 2018). We show that the kernels (and thus features) defined by the full parameter set and our restricted parameter set based on a few rows or columns are very similar in their numerical values and classification performance.
@@ 4 Experiments
We evaluate the effectiveness of the proposed RoCoFT method across various NLP tasks, including the General Language Understanding Evaluation (GLUE) benchmark, question answering, text summarization, common sense reasoning, and mathematical reasoning.
Baselines: For our baseline comparisons, we utilize prominent PEFT methods such as Adapter (Houlsby et al., 2019), Prompt Tuning (Lester et al., 2021), Prefix-Tuning (Li & Liang, 2021), (IA)3 (Liu et al., 2022), Bitfit (Zaken et al., 2021), LoRA (Hu et al., 2021), AdaLoRA (Zhang et al., 2023a), MAM Adapter (He et al., 2021), PROPETL (Zeng et al., 2023), LoKr (Edalati et al., 2022), (Wu et al., 2024) and LoHa (Hyeon-Woo et al., 2021). The experimental setup for the GLUE benchmark follows Xu et al. (2023), while question answering and text summarization tasks are conducted according to Zhang et al. (2023a).
@@
For LLM performance using RoCoFT, we conduct an extensive evaluation across thirteen benchmark datasets, covering both common sense reasoning and mathematical reasoning tasks, utilizing four LLMs: Bloom 7B (Le Scao et al., 2023), GPT-J 6B (Wang, 2021), LLaMa2-7B and LLaMA2-13B from Touvron et al. (2023). For common sense reasoning, we employ a wide range of datasets, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-easy and ARC-challenge (Clark et al., 2018), and OBQA (Mihaylov et al., 2018), ensuring comprehensive coverage of the model’s ability to handle diverse aspects of common sense reasoning. For mathematical reasoning, we use several specialized datasets, including MultiArith (Roy & Roth, 2016), GSM8K (Cobbe et al., 2021), AddSub (Hosseini et al., 2014), SingleEq (Koncel-Kedziorski et al., 2015), and SVAMP (Patel et al., 2021), to assess the model’s performance on arithmetic reasoning tasks. Detailed hyperparameter settings are provided in Appendix B. The implementation, environment setup, and hardware details of the experiments are given in Appendix C.
@@
RoCoFT3-Column 0.418M 62.95 94.69 89.18/90.94 90.85/90.45 89.86/86.38 86.76 91.89 79.21 85.55/90.69
RobertaLarge FT 355.3M 65.78 95.50 92.22/94.28 91.74/91.96 90.83/88.68 89.21 93.19 81.40 87.48/91.64
AdapterS 19.77M 65.33 96.37 89.88/90.23 92.58/92.42 91.19/87.11 91.00 94.31 85.25 88.24/90.05
Prompt-tuning 1.07M 61.13 94.61 73.04/76.29 78.51/78.99 80.74/75.16 68.15 89.13 60.29 75.70/76.81
Prefix-tuning 2.03M 59.01 95.76 88.24/89.37 90.92/91.07 88.88/85.45 89.30 93.32 74.01 84.93/88.63
(IA)3 1.22M 61.15 94.61 86.45/87.53 92.22/86.25 89.45/86.25 88.63 94.25 81.23 85.99/86.67
Bitfit 0.222M 67.01 96.10 90.93/92.13 91.93/93.38 89.48/86.43 89.98 94.47 87.73 88.45/90.14
LoRA 1.84M 64.47 96.67 87.50/88.19 91.66/91.44 90.15/86.91 90.76 95.00 79.78 86.99/88.85
AdaLoRA 2.23M 65.85 94.95 89.46/90.34 92.05/91.80 89.60/86.30 90.36 94.62 77.98 86.86/89.48
MAM Adapter 4.20M 67.39 95.81 90.12/92.07 92.44/92.18 90.87/86.65 90.62 94.31 86.62 88.52/90.30
PROPETL Adapter 5.40M 65.55 96.27 89.71/91.15 91.92/91.67 90.67/87.74 91.37 94.80 87.69 88.69/90.18
PROPETL Prefix 26.85M 62.24 96.17 90.04/91.92 90.70/90.49 89.30/86.30 90.33 94.73 79.71 86.65/89.57
PROPETL LoRA 4.19M 61.90 95.93 87.31/89.87 91.66/91.38 90.93/88.05 90.53 94.93 83.57 87.31/90.54
MoSLoRA 3.23M 67.27 96.17 89.96/92.67 90.97/91.72 90.12/87.68 90.29 94.73 82.41 87.43/90.69
RoCoFT1-Row 0.222M 65.70 96.63 89.97/90.79 91.81/92.07 90.17/86.15 90.73 94.20 85.31 88.07/89.67
RoCoFT3-Row 0.666M 67.39 96.69 91.05/92.19 92.10/92.10 90.82/86.11 90.98 94.85 87.83 88.96/90.13
RoCoFT1-Column 0.222M 64.89 96.60 89.12/90.24 91.96/92.10 90.17/85.83 90.81 94.17 85.71 87.92/89.39
RoCoFT3-Column 0.666M 67.18 96.67 89.88/91.47 92.52/92.31 91.38/87.12 91.13 94.85 87.82 88.93/90.30

### v3

@@ Abstract
We propose Row-Column Fine-Tuning (RoCoFT), a parameter-efficient finetuning method for large language models based on updating only a few rows and columns of the weight matrices in transformers. Through extensive experiments with medium size language models like RoBERTa and DeBERTa, and large language models (LLMs) liken Bloom-7B, Llama2-7B and Llama2-13B, we show that our method gives comparable accuracies to the state-of-the-art Parameter-Efficient Finetuning methods while also being more memory and computation-efficient. We also study the reason behind the effectiveness of our method with tools from Neural Tangent Kernel (NTK) theory. We empirically demonstrate that our kernel, constructed using a restricted set of row and column parameters, is numerically close to the full-parameter kernel and gives comparable classification performance. Ablation studies are conducted to investigate the impact of different algorithmic choices, including the robustness of RoCoFT to any selection of rows and columns, as well as the optimal rank for the effective implementation of our method.
RoCoFT: Efficient Finetuning of Large Language Models with Row-Column Updates
[…]
@@ 1 Introduction
Adapting Large Language Models (LLMs) to different downstream applications is the current prevailing paradigm for solving many Natural Language Processing (NLP) problems, such as sentiment analysis, machine translation, question answering, named entity recognition, and text summarization. LLMs like GPT-4 (Achiam et al., 2023) and Llama (Touvron et al., 2023) are trained on massive amounts of text data and contain billions of parameters. They give state-of-the-art performance on many NLP, mathematical reasoning (Hendrycks et al., 2021; Cobbe et al., 2021), and programming benchmarks (Jiang et al., 2024). Early works on transfer learning with pretrained LLMs, such as BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), use full finetuning, which updates all the parameters of the LLMs when adapting to downstream tasks. This approach becomes impractical as language models continue to scale up (Hoffmann et al., 2022), since a separate copy of the model parameters needs to be stored for each downstream application. Updating all the parameters is also prone to overfitting and the loss of LLM capabilities due to catastrophic forgetting (Kirkpatrick et al., 2017). Adaptor methods (Houlsby et al., 2019; Kowsher et al., 2024) solve this problem of finetuning LLMs by introducing extra modules called adaptors with a small set of independent parameters. Only the parameters in the adaptors need to be optimized during finetuning, and their small size makes it efficient to adapt an LLM to many different tasks. Parameter-Efficient Finetuning (PEFT) is the study of adapting LLMs to downstream applications by finetuning only a very small set of parameters. Many PEFT methods have been proposed, including the popular LoRA (Hu et al., 2021) and its variants (Zhang et al., 2023c; Edalati et al., 2022; Hyeon-Woo et al., 2021), prefix and prompt tuning (Li and Liang, 2021; Lester et al., 2021), and many other more advanced and complex adaptor methods (He et al., 2021; Zeng et al., 2023). These PEFT methods are effective in reducing the number of parameters required to adapt to downstream tasks, while maintaining performance close to full finetuning.
@@
Despite the numerous PEFT methods available, we pose a critical question: can we design even simpler PEFT methods capable of adapting LLMs to diverse downstream tasks in a more efficient way? A simpler method could not only enhance computational and storage efficiency but also offer deeper insights into why PEFT methods succeed as simpler methods are easier to analyze. We answer this question by presenting a new method called RoCoFT, where the LLMs can be efficiently adapted by updating only a small subset of rows or columns in the transformer block weight matrices.
We evaluate the effectiveness of our approach across several benchmarks on different language models. Our experimental results demonstrate that RoCoFT outperforms current PEFT techniques in accuracies, requires fewer trainable parameters, and has faster training times. We further analyze our method using Neural Tangent Kernel (NTK) theory (Jacot et al., 2018; Malladi et al., 2023), demonstrating that, for a pretrained LLM, the NTKs derived from a restricted set of rows and columns closely resemble those computed from the full parameter set. This substantiates the effectiveness of our proposed method and further suggests that most of the critical features for fine-tuning are already acquired during the pretraining phase.
@@
Our contributions are summarized as follows: (i) We introduce a new PEFT method called RoCoFT which gives comparable accuracies to state-of-the-art PEFT methods, while being more efficient in terms of memory and time complexity.These claims are validated through extensive experiments on language models of different sizes and many benchmark datasets. (ii) We analyze our method with empirical neural tangent kernels and show that these kernels are close to NTKs defined on the full parameter set, and they give comparable accuracies on many tasks when trained with kernel logistic regression. This explains why our method has performance close to full finetuning from the view of kernel methods. (iii) We perform extensive experiments and ablation studies on the design choices such as which and how many rows and columns to select to facilitate the implementation of our method.
@@ 2 Related Works
PEFT Methods: Parameter-Efficient Finetuning (PEFT) methods aim to finetune only a small number of existing or extra parameters of the LLM to achieve results comparable to finetuning all the parameters. Recently, numerous PEFT approaches have been proposed to advance this strategy. LoRA (Hu et al., 2021) and related methods (Zhang et al., 2023c; Kopiczko et al., 2023; Dettmers et al., 2024) modify existing weight matrices of the model by introducing trainable low-rank decomposition matrices, as adapters, into each layer of the Transformer (Vaswani et al., 2017) architecture. IA3 (Liu et al., 2022) is another adaptor method that only trains scaling vectors for the key, value, and feed-forward weight matrices in the attention mechanism for task adaptation. Prefix-Tuning (Li and Liang, 2021) and Prompt-Tuning (Lester et al., 2021; Kowsher et al., 2023) work by adding task-specific continuous vectors as contexts for inputs and only updates those parameters while keeping the original LLM parameters frozen. MAM adaptors (He et al., 2021) generalize from both LoRA and prefix-tuning under a unified framework. Our method is closer to LoRA and IA3 in that we modify the weight matrices in the transformer architecture. However, unlike these approaches, we introduce no extra parameters and modify the existing parameters in place. BitFit (Zaken et al., 2021) and LayerNorm Tuning (Zhao et al., 2023) finetune only the bias parameters and layernorm parameters respectively and are extremely parameter-efficient. However, unlike LoRA and our method they cannot increase the capacity of the finetuning model by increasing the rank, since the number of bias and layernorm parameters are fixed in a model.
Apart from low-rank adaptor methods Sparse Fine-Tuning is another group of PEFT methods that focuses on directly training only a very small subset of model parameters during finetuning. Sparsity can be achieved in two different ways, either by pruning after full finetuning or by selecting a sparse set of masks to train before finetuning. Diff pruning (Guo et al., 2021) encourages sparsity by using $l_{0}$ norm regularization during finetuning, while Ansell et al. (2022) makes use of the Lottery Ticket Hypothesis (Frankle and Carbin, 2018) to prune the weights after full finetuning. Unlike our proposed method they both require computational costs close to full finetuning. He et al. (2024) selects submatrix blocks as masks using maximal gradient change during warmup as criterion, while Sung et al. (2021) selects masks based on Fisher information. Both require some precomputation before a sparse mask can be selected for finetuning. Our method can be seen as belonging to both low-rank adaptor methods and sparse fine-tuning, as with few rows or columns chosen the updates are naturally both low-rank and sparse.
@@
Neural Tangent Kernels:  Jacot et al. (2018) and related studies (Lee et al., 2019) show that the training dynamics of an infinite-width multi-layer neural network with suitable Gaussian initialization can be completely described by a fixed kernel called the Neural Tangent Kernel (NTK). This result is further expanded in Yang (2020) to any architecture for which forward and backpropagation can be expressed via nonlinearities and matrix multiplications. Although these results are asymptotic, this interesting connection between deep neural networks and kernel methods allows many questions about neural networks to be studied via kernels. For example, Wei et al. (2022) studied the generalization error of representations learned by deep neural networks through kernel regression with NTKs. Recently Malladi et al. (2023) proposed to study the effect of finetuning LLMs through their NTKs, and provided theoretical support to their approach. In this paper, we continue along this line of work to use NTKs to analyze PEFT methods.
@@
Figure 1: A simplified overview of various PEFT methods and RoCoFT. Snowflake icon indicates frozen parameters while fire icon indicates trainable parameters.
@@ 3 RoCoFT
PEFT is a collection of methods for transferring pretrained LLMs to downstream tasks by optimizing only a very small set of (additional) parameters. Since most modern LLMs are based on the transformer architecture (Vaswani et al., 2017), there is a line of work in PEFT focusing on modifying the transformer block by freezing most of its parameters and training only a few additional parameters. The method proposed in Houlsby et al. (2019) adds adaptive layers to the transformer blocks, and only parameters in those adaptive layers need to be trained for effective transfer learning. LoRA (Hu et al., 2021) removes the need for adding adaptive layers by directly modifying the weight matrices used in the transformer blocks. There are multiple linear weight matrices in the transformer block taking up most of the parameters, including $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ for the query, key and value matrices in the attention mechanism, and also the weights $\mathbf{W}_{ff}$ for the MLP projection layers. LoRA makes use of a low-rank modification of these weight matrices
$\mathbf{W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A},$ (1)
where $\mathbf{W}_{0}$ is the pretrained weight matrix, $\mathbf{B}$ and $\mathbf{A}$ are low rank matrices of rank $r$, and $\mathbf{W}$ is the weight matrix after finetuning. In this formulation only $\mathbf{B}$ and $\mathbf{A}$ are updated. If $\mathbf{W}$ is of dimensions $d\times k$, $\mathbf{B}$ and $\mathbf{A}$ will be of dimensions $d\times r$ and $r\times k$ respectively. If $r\ll d,k$, this can lead to significant savings in terms of memory and runtime. IA3 in (Liu et al., 2022) also modifies the weight matrices in the transformer block, but instead of a low-rank modification they rescale the key and value matrices, and also the MLP layers using three learned vectors $\mathbf{l}_{k}$, $\mathbf{l}_{v}$ and $\mathbf{l}_{ff}$ dedicated to key, value and feedforward layers.
The success of these PEFT methods leads us to ask if there are even simpler methods for modifying the transformer block for effective finetuning. Here, we propose modifying only a few rows or columns in the weight matrices of the transformer block, for query, key, value weight matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ and also the weight matrices in the feedforward layer $\mathbf{W}_{ff}$. These can be expressed as
$\mathbf{W}=\mathbf{W}_{0}+\mathbf{R}\qquad\mbox{and}\qquad\mathbf{W}=\mathbf{W}_{0}+\mathbf{C},$ (2)
@@ 4 Experiments
We evaluate the effectiveness of the proposed RoCoFT method across various NLP tasks, including the General Language Understanding Evaluation (GLUE) benchmark, question answering, text summarization, common sense reasoning, and mathematical reasoning. We select rows or columns from the beginning of the order. This reduces the complexity of the selection process, as different selection strategies do not significantly affect performance. In other words, any row or column yields similar results, contributing to robustness. Further details are provided in Appendix 3.
Baselines: For our baseline comparisons, we utilize prominent PEFT methods such as Adapter (Houlsby et al., 2019), Prompt Tuning (Lester et al., 2021), Prefix-Tuning (Li and Liang, 2021), (IA)3 (Liu et al., 2022), Bitfit (Zaken et al., 2021), LoRA (Hu et al., 2021), AdaLoRA (Zhang et al., 2023b), MAM Adapter (He et al., 2021), PROPETL (Zeng et al., 2023), LoKr (Edalati et al., 2022), (Wu et al., 2024), SFT (Ansell et al., 2024), Diff Pruning (Guo et al., 2020), LoRAFA (Zhang et al., 2023a), Vera (Kopiczko et al., 2023), LoRA-XS (Bałazy et al., 2024) and LoHa (Hyeon-Woo et al., 2021). The experimental setup for the GLUE benchmark follows Xu et al. (2023), while question answering and text summarization tasks are conducted according to Zhang et al. (2023b).
@@
For LLM performance using RoCoFT, we conduct an extensive evaluation across thirteen benchmark datasets, covering both common sense reasoning and mathematical reasoning tasks, utilizing four LLMs: Bloom 7B (Le Scao et al., 2023), GPT-J 6B (Wang, 2021), LLaMa2-7B and LLaMA2-13B from Touvron et al. (2023). For common sense reasoning, we employ a wide range of datasets, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-easy and ARC-challenge (Clark et al., 2018), and OBQA (Mihaylov et al., 2018), ensuring comprehensive coverage of the model’s ability to handle diverse aspects of common sense reasoning. For mathematical reasoning, we use several specialized datasets, including MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AddSub (Hosseini et al., 2014), SingleEq (Koncel-Kedziorski et al., 2015), and SVAMP (Patel et al., 2021), to assess the model’s performance on arithmetic reasoning tasks. Detailed hyperparameter settings are provided in Appendix 10. The implementation, environment setup, and hardware details of the experiments are given in Appendix F.
@@
RoCoFT3-Column 0.249M 62.95 94.69 89.18/90.94 90.85/90.45 89.86/86.38 86.76 91.89 79.21 85.55/90.69